Find the zeros of the polynomial
f(x)= x^3+x^2-20x

Answer by LinnW(1048)

0= x^3+x^2-20x
0= x(x^2+x-20)
0= x(x+5)(x-4)
So x = 0, or x = -5 or x = 4